OpenStudy (anonymous):

Find the length of a rectangle with area of 16 sq. cm. and width of 8 cm. cm

OpenStudy (iurxos):

L*B=16 sq cm, and B=8, so L=16/8=2

OpenStudy (iurxos):

try to applie that to the other problems yourself ;) You can do it!